OpenStudy (anonymous):

Find the exact values of sin θ, cos θ, and tan θ if the terminal side of θ in standard position contains the given point. p(-3,0)

OpenStudy (dumbcow):

well normally you would draw a triangle but in this case you have a straight line from origin to (-3,0) sin = 0/3 = 0 cos = -3/3 = -1 tan = 0/-3 = 0